OpenStudy (anonymous):

Topic: Base e and Natural Logarithms Directions: Solve each equation or inequality round to nearest ten thousandth Problem: 1-2e^2x=-19

OpenStudy (anonymous):

\[-2 e ^{2x}=-19-1=-20,e ^{2x}=10,\ln e ^{2x}=\ln 10,2 x=\frac{ {\ln 10} }{ \ln e }=\ln 10\] x=?

OpenStudy (anonymous):

@surjithayer I am confused can you clear that confusion up please